# Backtracking Bipolar Magnetic Regions to their emergence: Two groups and their implication in the tilt measurements

### v1

@@ 4.2 Tilt angle of non-emerging BMRs
We compute the tilt angle $\gamma$ relative to the equator employing the usual definition of Y. M. Wang & N. R. Sheeley (1989). In this, BMRs located in the northern hemisphere and exhibiting Hale–Joy orientations are assigned positive tilt values, and the tilt value is kept within $\pm 90^{\circ}$. To get better statistical measure, we also combine the data of two hemispheres, by reversing the sign of the tilts of southern hemisphere.
@@
Figure 4(a, c, e) shows the tilt distribution of the BMRs that fall into the “non-emerging” groups. Across all three phases of their evolution — $T_{e}$, $T_{0}$, and $T_{m}$ (the matured phase of BMR)— the fitted Gaussian mean stays essentially at $0^{\circ}$, with no clear sign of systematic tilt. When we look at their Joy’s law behavior (Figure 4b, d, f), the picture becomes even clearer: at $T_{e}$ we see a scattered and even negative $\gamma_{0}$, which gradually shifts toward nearly $0^{\circ}$ by $T_{m}$, but without any meaningful dependence on latitude. In other words, their tilt evolution is completely inconsistent with the expected Joy’s law trend.
This behaviour of the tilt angle of non-emerging BMRs provides an independent check that this group of BMRs are not genuine emerging BMRs. Therefore, we need to isolate this group from the whole data set of BMRs identified in our detection code.
4.3 Joy’s law from the combined dataset
@@
Since non-emerging BMRs do not exhibit any characteristic ‘BMR-like’ behaviour, in this section we examine how combining the non-emerging and emerging classes affects the latitude dependence of tilt. Figure 5(a, c, e) shows the tilt distributions of BMRs at $T_{e}$, $T_{0}$, and $T_{m}$ (the matured phase) respectively. At $T_{e}$, the fitted Gaussian mean falls at $5.6^{\circ}\pm 1.7^{\circ}$, but with a very wide spread. Over time, the spread (quantified by $\sigma$) decreases significantly while the Gaussian mean does not change much as BMRs fully emerge and reach their peak flux. We note that Figure 5(e) should be the same as Figure 2(a) of A. Sreedevi et al. (2024). However, the smaller value of the Gaussian mean in the present data is because only a subset of BMRs could successfully be backtracked in the present study. Many high-flux BMRs that live for longer periods are not backtracked (as they either emerged near the east limb or on the far side of the Sun).
@@
Here in Figure 5(b, d, f) we also observe a Joy’s law trend, starting from $T_{e}$ and the noise around Joy’s law decreases steadily as BMRs evolve. However, here the Joy’s law trend is quite weaker than we found in our previous study (A. Sreedevi et al., 2025) excluding the non-emerging BMRs ($\gamma_{0}$ at $T_{e}$ was $25.98\pm 7.32$), demonstrating the influence of the non-emerging BMRs on the true signature of the BMR.
@@
In A. Sreedevi et al. (2025), we show that by separating the non-emerging BMRs, we find a robust Joy’s law at the time of emergence on the solar surface, consistent with the thin flux tube model, which suggests that the Coriolis force causes the tilt of emerging BMRs. If we do not isolate the non-emerging BMRs from the main data set as done here, then the scatter around Joy’s law is considerably increased, and in a limited data set, Joy’s law may not even be detected. This could be the case in previous studies (A. G. Kosovichev & J. O. Stenflo, 2008; H. Schunker et al., 2020) where no evidence of Joy’s law was found at the time of emergence of BMRs. Therefore, in any analysis of emerging BMRs, we need to remove the non-emerging BMRs carefully; otherwise, they may mask the actual properties of real BMRs that the Sun truly produces.
@@
The considerable amount of scatter in the tilt around Joy’s law observed at the time of first detection ($T_{e}$) originates from two sources. One is due to the inability to correctly detect the very early signature of the BMR and/or incorrect identification of the flux-weighted centroids of two poles, leading to uncertainties in tilt value. This is evident in Figure 7(a–d), where a BMR is wrongly identified [^1] By manually observing several BMRs we find that this type of wrong detection during back tracking phase is quite infrequent. during the backtracking phase of our automated code and produces incorrect tilts at $t<T_{e}$. Another source of scatter comes from a physical origin, and it is the turbulent convection. As explained in D. Longcope & A. R. Choudhuri (2002), during the rise of the BMR-forming flux tubes in the convection zone, they are buffeted by convective turbulence in the near-surface layer, and thus their tilts are easily disturbed. However, in the later phase, the Coriolis perturbation overpowers the short-time-scale perturbations of turbulence, which relax more rapidly. Thus, the tilt scatter reduces over the growth of the BMR as observed in our data. The decrease of the scatter in the tilt angle during the growth phase of BMR was also reported in the previous study of H. Schunker et al. (2020) using 153 emerging active regions.
5 Conclusion
@@
In this work, we have introduced an extension module to the Automatic Tracking Algorithm for BMRs (AutoTAB), designed specifically to capture the earliest signatures of BMR emergence. AutoTAB itself is a state-of-the-art tool that tracks the full near-side evolution of BMRs, producing a comprehensive catalog of their properties over time. However, existing tracking approaches—whether automated or reliant on human intervention—fall short in identifying the very onset of flux emergence. To address this, we developed a backtracking algorithm that takes the first detection time ($T_{0}$) from AutoTAB and traces the BMRs backward to the true moment of polarity appearance, thereby defining a new emergence time ($T_{e}$). Applying this method to the 12,000 unique BMRs in the AutoTAB catalog, we were able to backtrack and reliably determine $T_{e}$ for 3080 cases. This enabled us to systematically investigate their tilt angle distributions and latitude dependence at the moment of true emergence.
@@
By evaluating the flux and footpoint separation evolutions between $T_{e}$ and $T_{0}$, complemented by manual observations of several BMRs, we identified two distinct BMR groups: emerging and non-emerging. The emerging BMRs show significant flux growth, while the non-emerging ones do not. We show that the non-emerging BMRs have almost zero mean tilt and do not show any latitudinal preference like Joy’s law, even in their mature phase. The emerging BMRs show nonzero mean tilt and a robust Joy’s law even at the time of emergence on the solar surface, suggesting that they are caused by the Coriolis force, as explained by the thin flux tube model (A. Sreedevi et al., 2025). If we do not isolate the non-emerging BMRs from the whole data set obtained from the automated code, then Joy’s law trend at the time of emergence weakens as a large scatter accompanies it.
@@
Taken together, these results highlight a crucial implication: including non-emerging BMRs in statistical studies of solar active regions can severely mask the intrinsic properties of genuine, flux-emerging BMRs. To uncover the actual physical signatures of BMR emergence and their role in shaping Joy’s law, separating these two populations is essential.
@@
B.B.K. acknowledges the financial support from the Indian Space Research Organisation (project no. ISRO/RES/RAC-S/IITBHU/2024-25). {contribution}B.B.K. gave the initial idea. A.S., under the guidance of B.B.K., B.K.J., and D.B. analyzed magnetogram data, developed algorithms for BMR back tracking, and prepared the figures. R.G. analyzed Intensity Continuum data. A.S. and B.B.K. led the manuscript writing while B.K.J. commented on the text. All authors contributed to the discussion and presentation of the results.

### v2

@@ IV.2 Tilt angle of discarded BMRs
We compute the tilt angle $\gamma$ relative to the equator employing the usual definition of WS89. In this, BMRs located in the northern hemisphere and exhibiting Hale–Joy orientations are assigned positive tilt values, and the tilt value is kept within $\pm 90^{\circ}$. To get better statistical measure, we also combine the data of two hemispheres, by reversing the sign of the tilts of southern hemisphere.
Figure 4(a, c, e) shows the tilt distribution of the BMRs that fall into the “discarded” groups. Across all three phases of their evolution — $T_{e}$, $T_{0}$, and $T_{m}$ (the matured phase of BMR)— the fitted Gaussian mean stays essentially at $0^{\circ}$, with no clear sign of systematic tilt. The anti-Hale BMRs are the significant contributors in this discarded group. They contribute approximately 26%, 25% and 20% at $T_{e}$, $T_{0}$, and $T_{m}$ respectively. These numbers in the case of the emerging BMR group are: 23%, 13%, and 11%, respectively. These anti-Hale (Joy) and Hale-Anti-Joy BMRs, both of which have negative tilts, hinder the development of the polar field, causing variations in the solar cycle and thus posing challenges in solar cycle prediction (BK2023; KKS24; Dey25).
@@
When we look at their Joy’s law behavior in Figure 4b, d, f, we observe again a distinct picture for the discarded BMRs. At $T_{e}$ we see a scattered and even negative $\gamma_{0}$, which gradually shifts toward nearly $0^{\circ}$ by $T_{m}$, but without any meaningful dependence on latitude. In other words, their tilt evolution is completely inconsistent with the expected Joy’s law trend. This behavior of the tilt angle of discarded BMRs provides an independent check that this group of BMRs are not genuine emerging BMRs. Therefore, we need to isolate this group from the whole data set of BMRs identified in our detection code.
IV.3 Joy’s law from the combined dataset
@@
Since discarded BMRs do not exhibit any characteristic ‘BMR-like’ behavior, in this section we examine how combining the discarded and emerging classes affects the latitude dependence of tilt. Figure 5(a, c, e) shows the tilt distributions of BMRs at $T_{e}$, $T_{0}$, and $T_{m}$ (the mature phase) respectively. At $T_{e}$, the fitted Gaussian mean falls at $5.6^{\circ}\pm 1.7^{\circ}$, but with a very wide spread. Over time, the spread (quantified by $\sigma$) decreases significantly while the Gaussian mean does not change much as BMRs fully emerge and reach their peak flux. We note that Figure 5(e) should be the same as Figure 2(a) of SJ2024. However, the smaller value of the Gaussian mean in the present data is because only a subset of BMRs could successfully be backtracked in the present study. Many high-flux BMRs that live for longer periods are not backtracked (as they either emerged near the east limb or on the far side of the Sun).
@@
Here in Figure 5(b, d, f) we also observe a Joy’s law trend, starting from $T_{e}$ and the noise around Joy’s law decreases steadily as BMRs evolve. However, here the Joy’s law trend is quite weaker than we found in our previous study (sreedevi25a) excluding the discarded BMRs ($\gamma_{0}$ at $T_{e}$ was $25.98\pm 7.32$), demonstrating the influence of the discarded BMRs on the true signature of the BMR. This also iterates the importance of proper identification of the tilts in the region and need of cross-validation of tilt in various datasets (B2015; qin25). More specifically, our discarded BMRs isolated simply by observing the flux evolution will be excluded using the cross-validation of tilt from white-light measurements as done in qin25 because these discarded BMRs are mostly invisible in white-light observations.
@@
In sreedevi25a, we show that by separating the discarded BMRs, we find a robust Joy’s law at the time of emergence on the solar surface, consistent with the thin flux tube model, which suggests that the Coriolis force causes the tilt of emerging BMRs. If we do not isolate the discarded BMRs from the main data set as done here, then the scatter around Joy’s law is considerably increased, and in a limited data set, Joy’s law may not even be detected. This could be the case in previous studies (SK2008) where no evidence of Joy’s law was found at the time of emergence of BMRs. Therefore, in any analysis of emerging BMRs, we need to remove the discarded BMRs carefully; otherwise, they may mask the actual properties of real BMRs that the Sun truly produces.
@@
The considerable amount of scatter in the tilt around Joy’s law observed at the time of first detection ($T_{e}$) originates from two sources. One is due to the inability to correctly detect the very early signature of the BMR and/or incorrect identification of the flux-weighted centroids of two poles, leading to uncertainties in tilt value. This is evident in Figure 7(a–d), where a BMR is wrongly identified [^1] By manually observing several BMRs we find that this type of wrong detection during back tracking phase is quite infrequent. during the backtracking phase of our automated code and produces incorrect tilts at $t<T_{e}$. Another source of scatter comes from a physical origin, and it is the turbulent convection. As explained in LC2002, during the rise of the BMR-forming flux tubes in the convection zone, they are buffeted by convective turbulence in the near-surface layer, and thus their tilts are easily disturbed. However, in the later phase, the Coriolis perturbation overpowers the short-time-scale perturbations of turbulence, which relax more rapidly. Thus, the tilt scatter reduces over the growth of the BMR as observed in our data. The decrease of the scatter in the tilt angle during the growth phase of BMR was also reported in the previous study of SB2020 using 153 emerging active regions.
V Conclusion
@@
Various groups have developed algorithms for tracking the BMRs from magnetograms on the near-side, which are used to study their origins and evolutions. AutoTAB is one such state-of-the-art tool. However, existing tracking approaches—whether automated or reliant on human intervention—fall short in identifying the very onset of flux emergence. To address this, we developed a backtracking algorithm that takes the first detection time ($T_{0}$) from AutoTAB and traces the BMRs backward to the true moment of polarity appearance, thereby defining a new emergence time ($T_{e}$). Applying this method to the 12,000 unique BMRs in the AutoTAB catalog, we were able to backtrack and reliably determine $T_{e}$ for 3080 cases. This enabled us to systematically investigate their tilt angle distributions and latitude dependence at the moment of true emergence.
@@
By evaluating the flux and footpoint separation evolutions between $T_{e}$ and $T_{0}$, complemented by manual observations of several BMRs, we identified two distinct BMR groups: emerging and discarded. The emerging BMRs exhibit significant flux growth, whereas the discarded ones do not show such considerable growth; instead, they consist of the decaying BMRs with nearby active regions, ephemeral regions, or fragmented bipolar flux originating from a parent active region. We show that the discarded BMRs have almost zero mean tilt and do not show any latitudinal preference like Joy’s law, even in their mature phase. The emerging BMRs show nonzero mean tilt and a robust Joy’s law even at the time of emergence on the solar surface, suggesting that they are caused by the Coriolis force, as explained by the thin flux tube model (sreedevi25a). If we do not isolate the discarded BMRs from the whole data set obtained from the automated code, then Joy’s law trend at the time of emergence weakens as a large scatter accompanies it.
@@
Taken together, these results highlight a crucial implication: including discarded BMRs in statistical studies of solar active regions can severely mask the intrinsic properties of genuine, flux-emerging BMRs. To uncover the actual physical signatures of BMR emergence and their role in shaping Joy’s law, it is essential to separate these two populations.
The authors would like to thank the anonymous referee for the comments that helped in improving the manuscript. B.B.K. acknowledges the financial support from the Indian Space Research Organisation (project no. ISRO/RES/RAC-S/IITBHU/2024-25) and the Anusandhan National Research Foundation (ANRF) through the MATRIC program (file no. MTR/2023/000670).
{contribution}
@@
B.B.K. gave the initial idea. A.S., under the guidance of B.B.K., B.K.J., and D.B. analyzed magnetogram data, developed algorithms for BMR back tracking, and prepared the figures. R.G. analyzed Intensity Continuum data. A.S. and B.B.K. led the manuscript writing while B.K.J. commented on the text. All authors contributed to the discussion and presentation of the results.
@@
Figure 6: The time evolution of one ‘emerging BMR’ and two ‘discarded BMRs’ whose magnetograms are shown in Figure 1(e–h) and Figure 2(a–d), (e–h) repectively but are now searched in corresponding (near simultaneous) intensity continuum.
Figure 7: The same as Figure 1 but for a different BMR for which the backtracking phase ($t<T_{0}$) is wrongly identified.